My fourth grade daughter came home with this math question and I couldn't figure it out, please help me to explain to her how it's done. Here is the question
The perimeter of a triangle is 30", the first side of the triangle is 1" longer than the second side, the second side is 7" longer than the third side, what is the length of each side of the triangle.
Thank you in advance,
Jamie

Answer by mananth(16949)

The perimeter of a triangle is 30", the first side of the triangle is 1" longer than the second side, the second side is 7" longer than the third side, what is the length of each side of the triangle.
...
let the third side be x "
second side is 7 " longer = x+7
First side is 1" greater than second side = (x+7)+1=x+8
..
Perimeter = sum of all sides
x+x+7 + x+8 = 30
x+x+x+15=30
3x+15=30
-15
3x=15
/3
x= 5". ----------- third side
second side = x+7 = 5+7 =12"
third side = x+8 = 5+8 = 13 "
...
CHECK
5+12+13=30"
...
m.ananth@hotmail.ca